3.3.14 Let X and Y be independent, with X ∼ Bernoulli(1/2) and Y ∼ N(0, 1). Let
Z = X +Y and W = X −Y. Compute Var(Z), Var(W), Cov(Z, W), and Corr(Z, W)
1
Expert's answer
2020-12-21T17:52:56-0500

Given that X is Bernoulli random variable with parameter \frac{1}{2}

That is, X ∼ Bernoulli "(\frac{1}{2})"

Then, E(X) = p

"= \frac{1}{2}"

and V(X) = pq

"= \frac{1}{2} \times \frac{1}{2} \\

= \frac{1}{4}"

We know that "V(X) = E(X^2) -(E(X))^2"

"=> \frac{1}{4} = E(X)^2-(\frac{1}{2})^2 \\

=> E(X)^2 = \frac{1}{4}+\frac{1}{4}"

"= \frac{1}{2}"

Also, Y is a standard normal random variable.

That is, Y∼N(0,1) => E(Y) =0 and V(Y)=1

"V(Y) = E(Y^2) - (E(Y))^2 \\

=> 1 = E(Y^2) - (0)^2 \\

=> E(Y^2) = 1+0 \\

=1"

Given that X and Y are independent variables.

Now, given that Z=X+Y

Then,

"E(Z) = E(X)+E(Y) \\

= \frac{1}{2}+0 \\

= \frac{1}{2} \\

V(Z) = V(X)+V(Y)+2Cov(X,Y) \\

=V(X)+V(Y)+2(0)"

since X and Y are independent, Cov (X,Y) = 0

"= \frac{1}{4}+1 \\

= \frac{5}{4}"

Also,

"W=X-Y \\

=> E(W)=E(X)-E(Y) \\

= \frac{1}{2}-0 \\

= \frac{1}{2} \\

V(W) =V(X)+V(Y)-2Cov(X,Y) \\

= V(X)+V(Y)-2(0)"

since X and Y are independent,

"Cov (X,Y) = 0 \\

= \frac{1}{4}+1 \\

= \frac{5}{4} \\

Cov(Z,W)=E(ZW)-E(Z)E(W) \\

=E((X+Y)(X-Y))-E(Z)E(W) \\

= E(X^2-Y^2)-\frac{1}{2} \times \frac{1}{2} \\

=E(X^2)-E(Y^2) - \frac{1}{4} \\

= \frac{1}{2}-1- \frac{1}{4} \\

= -0.75"

"Corr(Z,W) = \frac{Cov(Z,W)}{\sqrt{V(Z)}\sqrt{V(W)}} \\

=\frac{-0.75}{\sqrt{\frac{5}{4}}\sqrt{\frac{5}{4}}} \\

= \frac{-0.75 \times 4}{5} \\

= -0.60"